OpenStudy (anonymous):

Factor the trinomial: x2 + 22x + 72 (x + 2)(x + 11) (x + 4)(x + 18) (x + 2)(x + 36) (x + 6)(x + 12)

OpenStudy (anonymous):

(x + 4)(x + 18)

OpenStudy (anonymous):

yeah, it's B - coz 4+18-22 and 4*18=72

OpenStudy (anonymous):

x2+18x+4x+72 x(x+18)+4(x+18) (x+4)(x+18)

OpenStudy (anonymous):

I meant: coz 4+18=22 and 4*18=72